# Dynamic Fusion Multimodal Network for SpeechWellness Detection

### v1

@@ 1 Introduction
Speech-based suicide risk monitoring offers several advantages, such as cost-effective and enabling continuous, non-invasive assessments [9]. Prior studies have demonstrated that individuals with suicidal tendencies often exhibit reduced speech efficiency, flattened prosody, monotonic delivery, and an overall lack of vocal energy [10]. In particular, spectral characteristics of speech can effectively reflect subtle psychomotor and emotional cues associated with suicidal ideation, such as variations in energy distribution, pitch, and harmonic content [11]. Moreover, acoustic features such as Mel-Frequency Cepstral Coefficients (MFCCs) and log-Mel spectrograms capture speech frequency content over time and have been linked to mental health conditions like depression. Kumar et al. showed that log-Mel spectrograms can effectively distinguish between depressed and non-depressed individuals [12]. Following this line, the 1st SpeechWellness detection challenge [^2] https://SpeechWellness.github.io was held with the goal of assessing suicide risk in adolescents through multimodal analysis of speech and text [13]. The challenge provided a unique dataset consisting of speech recordings from 600 Chinese teenagers aged 10 to 18, with 50% of the participants identified as at risk of suicide based on psychological scale assessments. The official baseline extracts audio features using Wav2vec 2.0 [14] and textual features using BERT [15]. However, the Wav2vec 2.0 in the audio branch extracts features directly from raw waveform inputs and lacks explicit frequency domain information, which can be highly informative for detecting suicide-related speech characteristics [16]. In addition, the total parameters of the challenge baseline system reach approximately 427 million, posing significant challenges for deployment on resource-constrained edge devices and limiting practical scalability.
@@ 2.2 Acoustic Branch in TF Domain
TF domain acoustic features like Mel-spectrogram and Mel-frequency cepstral coefficients (MFCC) are widely used in related research [21]. The Mel-spectrogram is computed by passing the audio signal through a bank of Mel-scale filters, producing features better aligned with human auditory perception. MFCC are derived by applying a logarithm to the Mel-filtered energies followed by a Discrete Cosine Transform (DCT), yielding compact low-dimensional vectors.

### v2

@@ 1 Introduction
Speech-based suicide risk monitoring offers several advantages, such as cost-effective and enabling continuous, non-invasive assessments [9]. Prior studies have demonstrated that individuals with suicidal tendencies often exhibit reduced speech efficiency, flattened prosody, monotonic delivery, and an overall lack of vocal energy [10]. In particular, spectral characteristics of speech can effectively reflect subtle psychomotor and emotional cues associated with suicidal ideation, such as variations in energy distribution, pitch, and harmonic content [11]. Moreover, acoustic features such as Mel-Frequency Cepstral Coefficients (MFCC) and Mel-spectrogram capture speech frequency content over time and have been linked to mental health conditions like depression. Kumar et al. showed that Mel-spectrogram can effectively distinguish between depressed and non-depressed individuals [12]. Following this line, the 1st SpeechWellness detection challenge [^2] https://SpeechWellness.github.io was held with the goal of assessing suicide risk in adolescents through multimodal analysis of speech and text [13]. The challenge provided a unique dataset consisting of speech recordings from 600 Chinese teenagers aged 10 to 18, with 50% of the participants identified as at risk of suicide based on psychological scale assessments. The official baseline extracts audio features using Wav2vec 2.0 [14] and textual features using BERT [15]. However, the Wav2vec 2.0 in the audio branch extracts features directly from raw waveform inputs and lacks explicit frequency domain information, which can be highly informative for detecting suicide-related speech characteristics [16]. In addition, the total parameters of the challenge baseline system reach approximately 427 million, posing significant challenges for deployment on resource-constrained edge devices and limiting practical scalability.
@@ 2.2 Acoustic Branch in TF Domain
TF domain acoustic features like Mel-spectrogram and MFCC are widely used in related research [21]. The Mel-spectrogram is computed by passing the audio signal through a bank of Mel-scale filters, producing features better aligned with human auditory perception. MFCC are derived by applying a Discrete Cosine Transform (DCT) to the Mel-filtered energies, yielding compact low-dimensional vectors.